OpenStudy (anonymous):

Solve for p. 3(p+q)=p A) q=-2/3p B)q=-3/2p c)p=-2/3q D)p=-3/2q I really need some help on this kind of question because I don't really understand it.

OpenStudy (anonymous):

3p+3q=p 2p=-3q p=-3q/2

OpenStudy (anonymous):

you need to multiply out on the left first, using the distributive law i.e. turn \(3(p+a)\) in to \(3p+3q\) then start with \[3p+3q=p\]

OpenStudy (anonymous):

you want now to get all the \(p\) stuff on one side of the equal sign, you can do that by subtracting \(3p\) from both sides, giving you \[3q=-2p\]

OpenStudy (anonymous):

final step to get \(p\) by itself is to divide both sides by \(-2\) to give a final answer of \[-\frac{3}{2}q=p\]

OpenStudy (anonymous):

Alright, thank you very much. I kinda get it now. @satellite73

OpenStudy (anonymous):

kinda is close enough i guess do a couple more and you will see it it is just like solving \[3(2+x)=x\] but with a \(q\) instead of a \(2\)

OpenStudy (anonymous):

yw how bout those pirates?!

OpenStudy (anonymous):

Yeah, I got to get better at this. I am good with equations that does not have so many variables in them. Anyways, the Pirates won last night. The score was 3-0, I think.

OpenStudy (anonymous):

(B.)